OpenStudy (anonymous):

write 3xxyyy using exponents?

OpenStudy (anonymous):

\[3x^2y^3\]

OpenStudy (anonymous):

You're just multiplying them.

OpenStudy (anonymous):

Thanks so much batman!

OpenStudy (doggy):

Basically just add the x's and the y's together

OpenStudy (anonymous):

thanks doggy!

OpenStudy (anonymous):

Not add...you multiply.

OpenStudy (anonymous):

oops thanks batman!

OpenStudy (anonymous):

3x2y3

OpenStudy (anonymous):

\[x \times x = x^2 , x + x = 2x\]

OpenStudy (anonymous):

thanks koehler493 and batman!